# Formal Degrees and Parabolic Induction: the Maximal Generic Case

### v2

@@ Abstract.
We study the compatibility of the formal degree conjecture and the parabolic induction process in the simplest nontrivial case for quasi-split $p$-adic groups. For a generic discrete series $\pi$ induced from an irreducible supercuspidal $\sigma$ of a maximal Levi subgroup, we compute the quotient $d(\pi)/d(\sigma)$ of formal degrees under some assumptions. As an application, we verify the conjecture for discrete series of split $\mathrm{G}_{2}$ supported on maximal Levi subgroups.
@@ 1. Introduction
The formal degree of a discrete series of $G$ is exactly its Plancherel measure in Harish-Chandra’s Plancherel formula. For real reductive groups, Harish-Chandra exhausted the discrete series and established the explicit Plancherel formula in 1970s. Following Harish-Chandra, Waldspurger completed the proof of the Plancherel formula for $p$-adic groups in [Wal03], but the Plancherel measure could not be explicitly determined since there’s no explicit classification of discrete series for $p$-adic groups. It was in [HII08a][HII08], assuming the local Langlands correspondence (conjectural for general $\mathbf{G}$), that an explicit formula for $d(\pi)$ was conjectured in terms of the adjoint $\gamma$-factor.
More precisely, let $\psi$ be a fixed non-trivial additive character of $F$ and $dg$ a specific Haar measure on $G$ depending on $\psi$ (cf. §3). Assuming the local Langlands correspondence for $\mathbf{G}$, let $(\varphi_{\pi},\rho_{\pi})$ be the refined Langlands parameter of $\pi$. Thus $\varphi_{\pi}:\mathrm{WD}_{F}\rightarrow{{}^{L}G}$ is an admissible homomorphism from the Weil-Deligne group $\mathrm{WD}_{F}=\mathrm{W}_{F}\times\mathrm{SL}_{2}(\mathbb{C})$ into the $L$-group ${{}^{L}G}$, and $\rho_{\pi}$ is an irreducible character of a certain component group (cf. §6.1). Let $\hat{G}$ (resp. $\hat{G}^{\natural}$) be the dual group of $\mathbf{G}$ (resp. $\mathbf{G}/\mathbf{A}$). Let $S_{\varphi_{\pi}}^{\natural}:=Z_{\hat{G}^{\natural}}(\varphi_{\pi}(\mathrm{WD}_{F}))$, and $\mathcal{S}_{\varphi_{\pi}}^{\natural}:=\pi_{0}(S_{\varphi_{\pi}}^{\natural})$ be the corresponding component group. It was conjectured that
@@
By general facts in representation theory of $p$-adic groups, every discrete series $\pi$ can be parabolically induced from some irreducible supercuspidal representation of a Levi subgroup. This means there exists an $F$-parabolic subgroup $\mathbf{P}$ with Levi decomposition $\mathbf{P}=\mathbf{M}\mathbf{N}$, a unitary supercuspidal representation $\sigma$ of $M$, and some $\lambda\in\mathfrak{a}^{*}_{M}$ such that $\pi$ is a subquotient of $i_{P}^{G}\sigma_{\lambda}$ (cf. §2). Then in order to reduce the formal degree conjecture to the supercuspidal case, it is natural to compute $\frac{d(\pi)}{d(\sigma)}$ and try to verify
@@
The main purpose of this paper is to study certain special cases of (1.1) using tools from local harmonic analysis. The ideas have already been used to deal with certain examples of classical groups in [GI14], while there are several technical difficulties for general cases:
@@
In this paper, we shall deal with the simplest nontrivial case. Let $\mathbf{G}$ be a quasi-split connected reductive group over $F$, with $\mathbf{P}=\mathbf{M}\mathbf{N}$ the standard maximal parabolic subgroup corresponding to $\theta=\Delta-\{\alpha\}$ for a positive simple root $\alpha\in\Delta$. Let $\tilde{\alpha}:=\langle\rho_{\mathbf{P}},\alpha^{\vee}\rangle^{-1}\rho_{\mathbf{P}}\in\mathfrak{a}_{\mathbf{M}}^{*}$ following Shahidi (cf. §4). Let $\pi$ be a generic discrete series representation of $G$ supported on $\mathbf{M}$, then by Shahidi’s classification there exists a unitary supercuspidal representation $\sigma$ of $M$ and a unique $j\in\{1,2\}$ such that $\pi$ is the unique generic subrepresentation of $i_{P}^{G}\sigma_{\tilde{\alpha}/j}$.
@@
  • $m$ is the index of $\mathrm{res}(X^{*}(\mathbf{M})^{\mathbf{G}})$ in $X^{*}(\mathbf{A}_{\mathbf{M}})^{\mathbf{G}}$ (cf. §4).
@@
This also shows how Langlands parameters could reflect information from local harmonic analysis and representation theory. Our computation is a combination of Heiermann’s formula for the quotient $d(\pi)/d(\sigma)$ (in terms of the residue of Harish-Chandra’s Plancherel density), and the Langlands-Shahidi method. The main difficulty is the explicit determination of some constants, which is somewhat technical but natural in principle. We also remark that for general non-generic representations, besides the invalidity of the Langlands-Shahidi method, the complexity of the behaviour of parameters also serves as an essential obstacle to the explicit computation of the adjoint $\gamma$-factor $\gamma(s,\pi,\mathrm{Ad},\psi)$ in this approach.
It remains to briefly explain the content of this paper. We first recollect some basic facts and fixes the notations in §2. In §3, we deal with the technical comparison of two different Haar measures. In §4 and §5, we make necessary preparations in local harmonic analysis, parameters, and compute the adjoint $\gamma$-factors. We then recall the formal degree conjecture and prove the main theorem in §6. Finally, as an application, we apply this result to discrete series of split $\mathrm{G}_{2}$ supported on a maximal Levi in §7.
@@ 2.1. General notations
  • $F$ is a non-archimedean local field of characteristic 00, with the ring of integers $\mathfrak{o}=\mathfrak{o}_{F}$ and a fixed uniformizer $\varpi\in\mathfrak{o}$. The order of the residue field $\mathfrak{o}/\varpi\mathfrak{o}$ is denoted by $q$.
@@
  • Denote by $\mathbf{T}_{0}$ the split component of $\mathbf{T}$, which is a maximal $F$-split torus of $\mathbf{G}$.
@@ 2.2. Harmonic analysis
In this subsection we recollect some basic notions and facts frequently used in local harmonic analysis. A standard reference is [Wal03]. Let $\mathbf{H}$ be an arbitrary connected reductive group over $F$ with $\mathbf{A}_{\mathbf{H}}$ the split component of its center.
@@
    be the group of unramified characters of $M$. We thus have a surjection $\mathfrak{a}_{\mathbf{H},\mathbb{C}}^{*}\twoheadrightarrow X^{\mathrm{ur}}(H)$ characterized by $\chi\otimes s\mapsto|\chi|^{s}$, which equipps $X^{\mathrm{ur}}(H)$ with a canonical complex structure. For $\lambda\in\mathfrak{a}_{\mathbf{H},\mathbb{C}}^{*}$, we denote by $\chi_{\lambda}\in X^{\mathrm{ur}}(H)$ its image.
  • To simplify the notations, given a smooth representation $\pi$ of $H=\mathbf{H}(F)$ and $\lambda\in\mathfrak{a}_{\mathbf{H},\mathbb{C}}^{*}$, we also denote by $\sigma_{\lambda}$ the representation $\sigma\otimes\chi_{\lambda}$.
@@
    $X_{0}^{\mathrm{ur}}(H):=\{\chi\in X^{\mathrm{ur}}(H):\chi(H)\subset\mathbb{C}^{1}\}\subset X^{\mathrm{ur}}(H)$
@@
    (Also cf. [Wal03, §I.1].)
  • Similarly, $\mathbf{A}_{\mathbf{G}}\subset\mathbf{A}_{\mathbf{M}}\subset\mathbf{M}\subset\mathbf{G}$ induces a canonical decomposition
@@
    It would be useful to identify $\mathfrak{a}_{\mathbf{M}}^{\mathbf{G}*}$ with $X^{*}(\mathbf{A}_{\mathbf{M}}/\mathbf{A}_{\mathbf{G}})\otimes\mathbb{R}$.
@@
  • For any closed subgroup $H$ of $G$, we equip $H$ with the Haar measure such that $K_{H}:=K_{G}\cap H$ has volume $1$. This is the standard choice of measures in local harmonic analysis.
  • Denote by $\bar{\mathbf{P}}=\mathbf{M}\bar{\mathbf{N}}$ the opposite parabolic of $\mathbf{P}$ (with respect to the choice of the maximal split torus $\mathbf{T}_{0}$) and $\delta_{\mathbf{P}}$ the modulus character of $\mathbf{P}$.
@@
We remark that this $\gamma(\mathbf{G}/\mathbf{M})$ is independent of the choice of $\mathbf{P}$ containing $\mathbf{M}$ (cf. [Wal03, pp. 240-241]), but does depend on the choice of $K_{G}$ (since the measure does).
@@ 2.3. Parameters and local factors
We recollect in this subsection some basic facts about Langlands parameters and local factors. For our purposes, stating them for quasi-split groups would suffice.
Let $\mathbf{G}$ be a quasi-split connected reductive group over $F$ with a fixed Borel subgroup $\mathbf{B}=\mathbf{T}\mathbf{U}$.
  • Denote by $\hat{G}$ the dual group of $\mathbf{G}$. Recall that $\hat{G}$ is then a connected reductive complex Lie group, whose root datum is the dual of the absolute root datum of $\mathbf{G}$.
  • After fixing a pinning of the based (absolute) root datum, $\Gamma:=\mathrm{Gal}(\overline{F}/F)$ acts on $\hat{G}$ via “pinned automorphisms”. The L-group of $\mathbf{G}$ is
    ${{}^{L}G}:=\hat{G}\rtimes\Gamma.$
    If $K/F$ is a finite Galois extension over which $G$ splits, then the action of $\Gamma$ factors through $\Gamma_{K/F}:=\mathrm{Gal}(K/F)$, and ${{}^{L}G}$ is essentially $\hat{G}\rtimes\Gamma_{K/F}$.
  • We fix a Borel subgroup $\hat{B}$ of $\hat{G}$ corresponding to the based root datum. A parabolic subgroup ${{}^{L}\!P}$ of ${{}^{L}G}$ is the normalizer $N_{{{}^{L}G}}(\hat{P})$ in ${{}^{L}G}$ of some parabolic $\hat{P}=\hat{M}\hat{N}$ of $\hat{G}$, with unipotent radical ${{}^{L}\!N}=\hat{N}$ and Levi ${{}^{L}\!M}=N_{{}^{L}\!P}(\hat{M})$. A parabolic ${{}^{L}\!P}$ of ${{}^{L}G}$ is said to be standard if ${{}^{L}\!P}\supset{{}^{L}\!B}$, which must be of the form ${{}^{L}\!P}=\hat{P}\rtimes\Gamma$, $\hat{P}\supset\hat{B}$ corresponding to a $\Gamma$-invariant subset of positive simple roots, with ${{}^{L}\!M}=\hat{M}\rtimes\Gamma$.
  • A standard parabolic subgroup $\mathbf{P}=\mathbf{M}\mathbf{N}$ of $\mathbf{G}$ then corresponds to a standard parabolic ${{}^{L}\!P}$ of ${{}^{L}G}$, which is relevant since $\mathbf{G}$ is assumed to be quasi-split. We could then identify the L-group of $\mathbf{M}$ with ${{}^{L}\!M}$. For more details, cf. [Bor79]. Following Shahidi’s convention, we denote by ${}^{L}\mathfrak{m}$ (resp. ${}^{L}\mathfrak{n}$) the Lie algebra of ${{}^{L}\!M}$ (resp. ${{}^{L}\!N}$).
@@
    • $\varphi:\mathrm{SL}_{2}(\mathbb{C})\rightarrow\hat{G}$ is algebraic;
    • $\varphi$ is continuous on $\mathrm{I}_{F}$ and $\varphi(\mathrm{Frob})\in\hat{G}$ is semisimple;
@@
  • Denote by $\Phi(\mathbf{G})$ (resp. $\Pi(\mathbf{G})$) the set of $\hat{G}$-conjugacy classes of L-parameters of $\mathbf{G}$ (resp. isomorphism classes of irreducible smooth representations of $G=\mathbf{G}(F)$).
@@
satisfying a series of natural properties, cf. [Bor79]. For any $\pi\in\Pi(\mathbf{G})$ (resp. $\varphi\in\Phi(\mathbf{G})$), $\mathrm{LL}(\pi):=\varphi_{\pi}$ (resp. $\mathrm{LL}^{-1}(\varphi):=\Pi_{\varphi}$) is called the L-parameter of $\pi$ (resp. the L-packet of $\varphi$).
Recall that we’ve fixed a nontrivial unitary character $\psi:F\rightarrow\mathbb{C}^{1}$. Assuming LLC and given a finite dimensional representation $r$ of the complex Lie group ${{}^{L}G}$, we denote by
@@
The choice of ${\rm rec}$ normalizes the LLC for split tori: for $\mathbf{G}=\mathrm{GL}_{1}$, a character $\chi:F^{\times}\rightarrow\mathbb{C}^{\times}$ corresponds to the parameter $\chi\circ\mathrm{rec}:\mathrm{W}_{F}\rightarrow\mathbb{C}^{\times}$. For example, the unramified character $\chi=|\cdot|^{\lambda}~{}(\lambda\in\mathbb{C})$ corresponds to the unramified character $\mathrm{W}_{F}\rightarrow\mathbb{C}^{\times}$ which maps $\mathrm{Frob}$ to $q^{-\lambda}$. We remark that in some literature such as [Hei06, §2.5], the authors use a normalization which gives $q^{\lambda}$, not $q^{-\lambda}$.

### v3

# references.bib

@@ Abstract.
We study the compatibility of the formal degree conjecture and the parabolic induction process in the simplest nontrivial case for quasi-split $p$-adic groups. For a generic discrete series $\pi$ induced from an irreducible supercuspidal $\sigma$ of a maximal Levi subgroup, we compute the quotient $d(\pi)/d(\sigma)$ of formal degrees with the assumption that the group is unramified. As an application, we verify the conjecture for discrete series of split $\mathrm{G}_{2}$ supported on maximal Levi subgroups.
@@ 1. Introduction
$\int_{G/A}(\pi(g)v,v^{\prime})\overline{(\pi(g)w,w^{\prime})}\mu(\dot{g})=\frac{1}{d(\pi)}(v,w)\overline{(v^{\prime},w^{\prime})},\quad v,v^{\prime},w,w^{\prime}\in V_{\pi}.$
The formal degree of a discrete series of $G$ is exactly its Plancherel measure in Harish-Chandra’s Plancherel formula. For real reductive groups, Harish-Chandra exhausted the discrete series and established the explicit Plancherel formula in 1970s. Following Harish-Chandra, Waldspurger completed the proof of the Plancherel formula for $p$-adic groups in [waldspurger-plancherel], but the Plancherel measure could not be explicitly determined since there’s no explicit classification of even supercuspidals for general $p$-adic groups. It was in [hiraga-ichino-ikeda][hiraga-ichino-ikeda-correction], assuming the local Langlands correspondence (conjectural for general $\mathbf{G}$), that an explicit formula for $d(\pi)$ was conjectured in terms of the adjoint $\gamma$-factor.
More precisely, let $\psi$ be a fixed non-trivial additive character of $F$ and $\mu_{\psi}(g)$ a specific Haar measure on $G$ depending on $\psi$ (cf. §3). Assuming the local Langlands correspondence for $\mathbf{G}$, let $(\varphi_{\pi},\rho_{\pi})$ be the refined Langlands parameter of $\pi$. Thus $\varphi_{\pi}:\mathrm{WD}_{F}\rightarrow{{}^{L}G}$ is an admissible homomorphism from the Weil-Deligne group $\mathrm{WD}_{F}=\mathrm{W}_{F}\times\mathrm{SL}_{2}(\mathbb{C})$ into the $L$-group ${{}^{L}G}$, and $\rho_{\pi}$ is an irreducible character of a certain component group (cf. §6.1). Let $\widehat{G}$ (resp. $\widehat{G}^{\natural}$) be the dual group of $\mathbf{G}$ (resp. $\mathbf{G}/\mathbf{A}$), so that $\widehat{G}^{\natural}$ is a subgroup of $\widehat{G}$. Let $S_{\varphi_{\pi}}^{\natural}:=Z_{\widehat{G}^{\natural}}(\varphi_{\pi}(\mathrm{WD}_{F}))$, and $\mathcal{S}_{\varphi_{\pi}}^{\natural}:=\pi_{0}(S_{\varphi_{\pi}}^{\natural})$ be the corresponding component group. It was conjectured that
@@
By general facts in representation theory of $p$-adic groups, every discrete series $\pi$ can be parabolically induced from some irreducible supercuspidal representation of a Levi subgroup. This means
  • there exists an $F$-parabolic subgroup $\mathbf{P}$ with Levi decomposition $\mathbf{P}=\mathbf{M}\mathbf{N}$, an irreducible unitary supercuspidal representation $\sigma$ of $M$, and some $\lambda\in\mathfrak{a}^{*}_{\mathbf{M}}=X^{*}(\mathbf{M})\otimes\mathbb{R}$ such that $\pi$ is a subquotient of $i_{P}^{G}\sigma_{\lambda}$ ($\sigma_{\lambda}:=\sigma\otimes\chi_{\lambda}$, where $\chi$ is the unramified character of $M$ associated to $\lambda\in\mathfrak{a}^{*}_{\mathbf{M}}$, cf. §2).
Then in order to reduce the formal degree conjecture to the supercuspidal case, it is natural to compute $\frac{d(\pi)}{d(\sigma)}$ and try to verify
@@
The main purpose of this paper is to study certain special cases of (1.1) using tools from local harmonic analysis. The ideas have already been used to deal with certain examples of classical groups in [gan-ichino-formal-degree], while there are several technical difficulties for general cases:
@@
In this paper, we shall deal with the simplest nontrivial case. Let $\mathbf{G}$ be a quasi-split connected reductive group over $F$ with fixed Borel subgroup, and $\mathbf{P}=\mathbf{M}\mathbf{N}$ standard maximal parabolic subgroup with $\tilde{\alpha}\in\mathfrak{a}_{\mathbf{M}}^{*}$ the corresponding fundamental weight (cf. §4 for their precise meanings).
Let $\pi$ be a generic discrete series representation of $G$ supported on $\mathbf{M}$, then by Shahidi’s classification there exists an irreducible unitary supercuspidal representation $\sigma$ of $M$ and a unique $j\in\{1,2\}$ such that $\pi$ is the unique generic subrepresentation of $i_{P}^{G}\sigma_{\tilde{\alpha}/j}$.
@@
  • $m$ is the index of $\mathrm{res}(X^{*}(\mathbf{M})^{\mathbf{G}})$ in $X^{*}(\mathbf{A}_{\mathbf{M}})^{\mathbf{G}}$ (cf. §4 for the notations).
@@
This also shows how Langlands parameters could reflect information from local harmonic analysis and representation theory. As an application, we apply this to prove the formal degree conjecture for discrete series of split $\mathrm{G}_{2}$ supported on a maximal Levi (based on the local Langlands for $\mathrm{G}_{2}$ established recently by Gan-Savin [gan-savin-g2theta][gan-savin-g2-llc]).
Our computation is a combination of Heiermann’s formula for the quotient $d(\pi)/d(\sigma)$ (in terms of the residue of Harish-Chandra’s Plancherel density), and the Langlands-Shahidi method. The main difficulty is the explicit determination of some constants, which is somewhat technical but natural in principle. We also remark that for general non-generic representations, besides the invalidity of the Langlands-Shahidi method, the complexity of the behaviour of parameters also serves as an essential obstacle to the explicit computation of the adjoint $\gamma$-factor $\gamma(s,\pi,\mathrm{Ad},\psi)$ in this approach.
It remains to briefly explain the content of this paper. We first recollect some basic facts and fix the notations in §2. In §3, we deal with the technical comparison of two different Haar measures. In §4 and §5, we make necessary preparations in local harmonic analysis, parameters, and compute the adjoint $\gamma$-factors. We then recall the formal degree conjecture and prove the main theorem in §6. Finally, as an application, we apply this result to discrete series of split $\mathrm{G}_{2}$ supported on a maximal Levi in §7.
@@ 2.1. General notations
  • $F$ is a non-archimedean local field of characteristic 00, with the ring of integers $\mathfrak{o}=\mathfrak{o}_{F}$, its maximal ideal $\mathfrak{p}=\mathfrak{p}_{F}$, and a fixed uniformizer $\varpi\in\mathfrak{p}$. The order of the residue field $\mathfrak{o}/\mathfrak{p}$ is denoted by $q$. We fix an algebraic closure $\overline{F}$ of $F$.
@@
  • Denote by $\mathbf{T}_{0}$ the split component of $\mathbf{T}$, which is a maximal $F$-split torus of $\mathbf{G}$, and $X_{*}(\mathbf{T}_{0})$ the set of cocharacters of $\mathbf{T}_{0}$.
@@ 2.2. Harmonic analysis
In this subsection we recollect some basic notions and facts frequently used in local harmonic analysis. A standard reference is [waldspurger-plancherel]. Let $\mathbf{H}$ be an arbitrary connected reductive group over $F$ with $\mathbf{A}_{\mathbf{H}}$ the split component of its center.
@@
    be the group of unramified characters of $H$. We thus have a surjection $\mathfrak{a}_{\mathbf{H},\mathbb{C}}^{*}\twoheadrightarrow X^{\mathrm{ur}}(H)$ characterized by $\chi\otimes s\mapsto|\chi|^{s}$, which equips $X^{\mathrm{ur}}(H)$ with a canonical complex structure. For $\lambda\in\mathfrak{a}_{\mathbf{H},\mathbb{C}}^{*}$, we denote by $\chi_{\lambda}\in X^{\mathrm{ur}}(H)$ its image.
  • To simplify the notations, given a smooth representation $\sigma$ of $H=\mathbf{H}(F)$ and $\lambda\in\mathfrak{a}_{\mathbf{H},\mathbb{C}}^{*}$, we also denote by $\sigma_{\lambda}$ the representation $\sigma\otimes\chi_{\lambda}$.
@@
    $X_{0}^{\mathrm{ur}}(H):=\{\chi\in X^{\mathrm{ur}}(H)\mid\chi(H)\subset\mathbb{C}^{1}\}\subset X^{\mathrm{ur}}(H)$
@@
    (Also cf. [waldspurger-plancherel, §I.1].) Similarly, $\mathbf{A}_{\mathbf{G}}\subset\mathbf{A}_{\mathbf{M}}\subset\mathbf{M}\subset\mathbf{G}$ induces a canonical decomposition
@@
    It would be convenient to identify $\mathfrak{a}_{\mathbf{M}}^{\mathbf{G}*}$ with $X^{*}(\mathbf{A}_{\mathbf{M}}/\mathbf{A}_{\mathbf{G}})\otimes\mathbb{R}$.
@@
  • For any closed subgroup $H$ of $G$, we equip $H$ with the Haar measure such that $K_{H}:=K_{G}\cap H$ has volume one. This is the standard choice of measures in local harmonic analysis.
  • Recall $\bar{\mathbf{P}}=\mathbf{M}\bar{\mathbf{N}}$ is the opposite parabolic of $\mathbf{P}$ (with respect to the choice of the maximal split torus $\mathbf{T}_{0}$ and a Borel $\mathbf{B}$). Let $\delta_{\mathbf{P}}$ be the modulus character of $\mathbf{P}$.
@@
We remark that this $\gamma(\mathbf{G}/\mathbf{M})$ is independent of the choice of $\mathbf{P}$ containing $\mathbf{M}$ (cf. [waldspurger-plancherel, pp. 240-241]), but does depend on the choice of $K_{G}$ (since the measure does).
@@ 2.3. Parameters and local factors
We recollect in this subsection some basic facts about Langlands parameters and local factors. For our purposes, stating them for quasi-split groups would suffice. Let $\mathbf{G}$ be a quasi-split connected reductive group over $F$ with a fixed Borel subgroup $\mathbf{B}=\mathbf{T}\mathbf{U}$.
  • Denote by $\widehat{G}$ the dual group of $\mathbf{G}$. Recall that $\widehat{G}$ is then a connected reductive complex Lie group, whose root datum is the dual of the absolute root datum of $\mathbf{G}$.
  • After fixing a pinning of the based (absolute) root datum, $\Gamma:=\mathrm{Gal}(\overline{F}/F)$ acts on $\widehat{G}$ via “pinned automorphisms”. The L-group of $\mathbf{G}$ is
    ${{}^{L}G}:=\widehat{G}\rtimes\Gamma.$
    If $K/F$ is a finite Galois extension over which $G$ splits, then the action of $\Gamma$ factors through $\Gamma_{K/F}:=\mathrm{Gal}(K/F)$, and ${{}^{L}G}$ is essentially $\widehat{G}\rtimes\Gamma_{K/F}$.
  • We fix a Borel subgroup $\widehat{B}$ of $\widehat{G}$ corresponding to the based root datum. A parabolic subgroup ${{}^{L}\!P}$ of ${{}^{L}G}$ is the normalizer $N_{{{}^{L}G}}(\widehat{P})$ in ${{}^{L}G}$ of some parabolic $\widehat{P}=\widehat{M}\widehat{N}$ of $\widehat{G}$, with unipotent radical ${{}^{L}\!N}=\widehat{N}$ and Levi ${{}^{L}\!M}=N_{{}^{L}\!P}(\widehat{M})$. A parabolic ${{}^{L}\!P}$ of ${{}^{L}G}$ is said to be standard if ${{}^{L}\!P}\supset{{}^{L}\!B}$, which must be of the form ${{}^{L}\!P}=\widehat{P}\rtimes\Gamma$, $\widehat{P}\supset\widehat{B}$ corresponding to a $\Gamma$-invariant subset of positive simple roots, with ${{}^{L}\!M}=\widehat{M}\rtimes\Gamma$.
  • A standard parabolic subgroup $\mathbf{P}=\mathbf{M}\mathbf{N}$ of $\mathbf{G}$ then corresponds to a standard parabolic ${{}^{L}\!P}$ of ${{}^{L}G}$, which is relevant since $\mathbf{G}$ is assumed to be quasi-split ([borel-automorphic-L-function, §3]). We could then identify the L-group of $\mathbf{M}$ with ${{}^{L}\!M}$. For more details, cf. [borel-automorphic-L-function]. We denote by $\widehat{\mathfrak{m}}$ (resp. $\widehat{\mathfrak{n}}$) the Lie algebra of ${\widehat{M}}$ (resp. ${\widehat{N}}$).
@@
    • $\varphi:\mathrm{SL}_{2}(\mathbb{C})\rightarrow\widehat{G}$ is algebraic;
    • $\varphi$ is continuous on $\mathrm{I}_{F}$ and $\varphi(\mathrm{Frob})\in\widehat{G}$ is semisimple;
@@
  • Denote by $\Phi(\mathbf{G})$ (resp. $\Pi(\mathbf{G})$) the set of $\widehat{G}$-conjugacy classes of L-parameters of $\mathbf{G}$ (resp. isomorphism classes of irreducible smooth representations of $G=\mathbf{G}(F)$).
@@
satisfying a series of properties. For any $\pi\in\Pi(\mathbf{G})$ (resp. $\varphi\in\Phi(\mathbf{G})$), $\mathrm{LL}(\pi):=\varphi_{\pi}$ (resp. $\mathrm{LL}^{-1}(\varphi):=\Pi_{\varphi}$) is called the L-parameter of $\pi$ (resp. the L-packet of $\varphi$).
Recall that we have fixed a nontrivial unitary character $\psi:F\rightarrow\mathbb{C}^{1}$. Assuming LLC and given a finite dimensional representation $r$ of the complex Lie group ${{}^{L}G}$, we denote by
@@
The choice of ${\rm rec}$ normalizes the LLC for split tori: for $\mathbf{G}=\mathrm{GL}_{1}$, a character $\chi:F^{\times}\rightarrow\mathbb{C}^{\times}$ corresponds to the parameter $\chi\circ\mathrm{rec}:\mathrm{W}_{F}\rightarrow\mathbb{C}^{\times}$. For example, the unramified character $\chi=|\cdot|^{\lambda}~{}(\lambda\in\mathbb{C})$ corresponds to the unramified character $\mathrm{W}_{F}\rightarrow\mathbb{C}^{\times}$ which maps $\mathrm{Frob}$ to $q^{-\lambda}$. We remark that in some literature such as [heiermann-orbites, §2.5], the authors use a normalization which gives $q^{\lambda}$, not $q^{-\lambda}$.